How many grams of CaCIz are needed to make 558.4 g
How many grams of CaCIz are needed to make 558.4 g
How many grams of CaCIz are needed to make 558.4 g
Answer #1

Q1.

mass needed for

m = 558.4 g of solution in 40.5% CaCl2

so..

%mm = mass of CaCl2 / total mass *100%

40.5/100 = mass of salt / (mass of salt + water)

0.405*MS + 0.405*W = MS

0.405*W = (1-0.405)*MS

W = 558.4-MS

0.405*(558.4-MS) = (1-0.405)*MS

0.405*(558.4) - 0.405MS = MS - 0.405*MS

MS(1-0.405+0.405) = 226.152

MS = (226.152)/((1-0.405+0.405) )

MS = 226.152 mass of salt required

b

for wter

W = 558.4-MS = 558.4-226.152 = 332.248 g of water

c)

x = 37.5%

Mtotal = 697.8 g, find CaCl2:

mass of CaCl2 = x*Mtotal = 37.5/100 * 697.8 = 261.675 g of CaCl2

mass of water = 697.8-261.675 = 436.125g

For all further questions, please post multiple questions in multiple set of Q&A. We ar enot allowed to answer to multiple questions in a single Q&A post.